Mariam uses a pottery wheel to make three bowls in 2 hours. Duane makes five bowls in 3 hours. Who makes bowls faster Mariam or Duane?
Found 2 solutions by oberobic, MathTherapy:
Answer by oberobic(2304)

Mariam's bowl-making speed = 1/M bowls/hr
1/M * 2hr = 3 bowls
.
Duane's bowl-making speed = 1/D bowls/hr
1/D * 3 hr = 5 bowls
.
1/M *2 = 3
2/M = 3
2 = 3M
3M = 2
M = 2/3 bowls/hr
.
1/D *3 = 5
3/D = 5
3 = 5D
5D = 3
D = 3/5 bowls/hr
.
2/3 = .67
3/5 = .60
.
So, Mariam makes bowls faster than Duane.

Answer by MathTherapy(10858)

Mariam uses a pottery wheel to make three bowls in 2 hours. Duane makes five bowls in 3 hours. Who makes bowls faster Mariam or Duane?

Mariam: 3 bowls in 2 hours, or 3/2 bowls, or 1.5 bowls, per hour
Duane: 5 bowls in 3 hours, or 5/3 bowls, or 1.6666 bowls, per hour

OR

At 1.5 bowls per hour, Mariam will make 4.5 bowls (3 * 1.5) in 3 hours. In that time (3 hours), Duane will have made 5 bowls.

Who’s faster? You decide!!

Send comments and “thank-yous” to “D” at MathMadEzy@aol.com